# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ Appendix C Models Inference and Training Details
?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen. […] Ich will ganz klar sagen: Das Boot ist nicht voll; das Boot ist niemals voll. Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […]. ? ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us. […] I want to state very clearly: the boat is not full; the boat is never full. Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
(16) Predicted label: group-based anti-solidarity towards refugees
@@
?[…] Wenn die Zahlen weiterhin auf diesem hohen Niveau bleiben, dann wird Deutschland über kurz oder lang überfordert sein. Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […]. allein im September sind mehrere zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen. Sie wurden nicht registriert. Sie wurden nicht kontrolliert. […] Damit besteht auch ein großes Sicherheitsrisiko. ? ?[…] If the numbers remain at this high level, Germany will sooner or later be overburdened. Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […]. In September alone, several tens of thousands of refugees and asylum seekers came to Germany. They were not registered. They were not controlled. […] This poses a major security risk.?
(17) Predicted label: exchange-based anti-solidarity towards refugees
@@
?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug nach derzeitigem Stand um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind. In Deutschland, meine Damen und Herren, lebt bereits eine große soziale Unterschicht […]. Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant. Meine Damen und Herren, Deutschland ist kein reiches Land, sondern ein leistungsfähiges. ? ?[…] each year about 140,000 highly qualified people leave this country, while in return, according to the current situation, about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate. In Germany, ladies and gentlemen, there already lives a large social underclass […]. Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people. Ladies and gentlemen, Germany is not a rich country, but a capable one.?
Table 11: Examples for the 2009-2025 period discussed in Section 4.3. Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(18) Predicted label: group-based anti-solidarity towards refugees
@@
(19) Predicted label: group-based anti-solidarity towards refugees
@@
(Dec 5, 2023)
?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt. […] Dafür werden bereits Zimmer in angeblich christlichen Altenheimen gekündigt, weil angebliche Flüchtlinge lukrativer sind, Messerzuwanderung inklusive. […] Bei diesem Recht auf Wohnungstausch der Linken handelt es sich um eine Mischung aus Neid zulasten alter Leute und Zwang zulasten von Vermietern. Es ist die Fortsetzung einer inländerfeindlichen… ? ?[…] elderly people are supposed to move out of their large apartments because green-left politics invites millions of guests from Africa and the Middle East while simultaneously stifling housing construction. […] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration [^12] Messereinwanderung: a propagandistic term coined by the AfD implying migrants bring knife crime. included. […] This ?right to apartment swaps? proposed by the Left is a mixture of envy at the expense of elderly people and coercion at the expense of landlords. It is the continuation of an anti-native policy…?
(20) Predicted label: exchange-based solidarity towards immigrants
@@
(21) Predicted label: group-based anti-solidarity towards migrants
@@
Table 12: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(22) Predicted label: group-based anti-solidarity towards immigrants
@@
(23) Predicted label: group-based anti-solidarity towards foreigners
Misclassified; true: group-based solidarity
@@
(Apr 10, 2024)
?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht. Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird. Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […]. Sie ?übersehen? in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben. Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen. Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist. […] ? ?Let us turn to the AfD’s draft and to what makes it simply atrocious. You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved. Not only does this place all foreigners under general suspicion […], but it also violates EU law […]. You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency. But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is. […]?
(24) Predicted label: group-based anti-solidarity towards migrants
Party: SPD
(Dec 9, 2024)
@@
(25) Predicted: group-based anti-solidarity towards migrants
Misclassified; true: empathic solidarity
Party: Die Linke
(June 5, 2025)
?[…] Migration wird pauschal als sicherheitspolitisches Problem dargestellt. Die AfD führt die sogenannte gescheiterte Migrationspolitik in einer Liste von Krisen auf – neben dem Krieg in der Ukraine, der Coronapandemie und der hybriden Bedrohung. Damit betreibt sie eine gefährliche und spaltende Gleichsetzung. Sie erklärt das bloße Vorhandensein von Migration zur Bedrohung der inneren und äußeren Sicherheit und nennt sie in einem Atemzug mit Terrorismus, Krieg und Katastrophen. ? ?Migration is presented across the board as a security policy problem. The AfD lists the so-called failed migration policy alongside the war in Ukraine, the coronavirus pandemic, and hybrid threats. In doing so, it engages in a dangerous and divisive equation. It declares the mere existence of migration to be a threat to internal and external security, mentioning it in the same breath as terrorism, war, and disasters.?
Table 13: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
@@
(26) Predicted label: exchange-based solidarity towards immigrants
@@
(27) Predicted label: empathic solidarity towards migrants
@@
Table 14: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.

### v2

@@ Appendix E Label Distribution Distance Metrics
?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen. […] Ich will ganz klar sagen: Das Boot ist nicht voll […]. Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […]. ? ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us. […] I want to state very clearly: the boat is not full […]. Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
(19) Dominant DSL subtype score: group-based anti-solidarity towards refugees
@@
?[…] Deutschland [wird] über kurz oder lang überfordert sein. Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […]. allein im September sind zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen. Sie wurden nicht registriert. Sie wurden nicht kontrolliert. […] Damit besteht auch ein Sicherheitsrisiko. ? ?[…] Germany will sooner or later be overburdened. Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […]. In September alone, several tens of thousands of refugees and asylum seekers came to Germany. They were not registered. They were not controlled. […] This poses a major security risk.?
Table 20: Examples for the 2009-2025 period discussed in Section˜4.3. Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(20) Dominant DSL subtype score: exchange-based anti-solidarity towards refugees
@@
?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug […] um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind. In Deutschland lebt bereits eine große soziale Unterschicht […]. Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant. […] Deutschland ist kein reiches Land, sondern ein leistungsfähiges. ? ?[…] each year about 140,000 highly qualified people leave this country, while in return, […] about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate. In Germany, there already lives a large social underclass […]. Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people. […] Germany is not a rich country, but a capable one.?
(21) Dominant DSL subtype score: group-based anti-solidarity towards refugees
@@
(22) Dominant DSL subtype score: group-based anti-solidarity towards refugees
@@
(12 May, 2023)
?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt. […] Dafür werden bereits Zimmer in angeblich christlichen Altenheimen gekündigt, weil angebliche Flüchtlinge lukrativer sind, Messerzuwanderung inklusive. […] Bei diesem Recht auf Wohnungstausch der Linken handelt es sich um eine Mischung aus Neid zulasten alter Leute und Zwang zulasten von Vermietern. Es ist die Fortsetzung einer inländerfeindlichen… ? ?[…] elderly people are expected to leave their large apartments because green-left politics invites millions of guests from Africa and the Middle East while choking off housing construction. […] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration included. […] This ?right to apartment swaps? proposed by the Left is a mixture of envy at the expense of the elderly and coercion at the expense of landlords. It is the continuation of an anti-native [policy]…?
(23) Dominant DSL subtype score: exchange-based solidarity towards immigrants
@@
Table 21: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(24) Dominant DSL subtype score: group-based anti-solidarity towards migrants
@@
(25) Dominant DSL subtype score: exchange-based anti-solidarity towards immigrants
@@
(26) Dominant DSL subtype score: group-based anti-solidarity towards migrants
@@
(12 Sept, 2024)
@@
Table 22: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
@@
(27) Dominant DSL subtype score: exchange-based solidarity towards immigrants
@@
(28) Dominant DSL subtype score: compassionate solidarity towards migrants
@@
(29) Dominant DSL subtype score: compassionate anti-solidarity towards asylum seekers
Party: CDU/CSU
(19 Feb 2016)
?Das sind aber nicht nur Flüchtlinge, die Anspruch auf Schutz haben. […] bei mehr als der Hälfte [konnte] keine Schutzbedürftigkeit festgestellt werden. Wenn wir bei Asylbewerbern, bei denen schon bei der Ankunft absehbar ist, dass sie keine Bleibeaussicht haben, nicht in kürzester Zeit entscheiden und sie in ihre Herkunftsländer zurückführen, werden wir noch mehr Anreize schaffen, ohne berechtigten Grund nach Deutschland zu kommen. […] Das ist nicht nur ein klares Signal im Sinne von ?Wir nehmen nur wirklich Schutzbedürftige auf?, sondern entlastet auch deutlich unsere Kommunen bei der Unterbringung. Man kann dabei auch nicht von einer Einschränkung oder Aushöhlung des Asylrechts sprechen. ? ?These are not only refugees who are entitled to protection. […] in more than half of the cases no need for protection could be established. If, in the case of asylum seekers for whom it is already apparent upon arrival that they have no prospect of remaining, we do not make a decision as quickly as possible and return them to their countries of origin, we will create even more incentives to come to Germany without justified reason. […] This is not only a clear signal in the sense of ?We only take in those who genuinely need protection,? but it also significantly relieves our municipalities with regard to accommodation. Nor can one speak of a restriction or hollowing-out of the right to asylum in this context.?
@@
Table 23: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Appendix F Annotation Process Details
Annotators first assigned a high-level stance label and, where applicable, an (anti-)solidarity subtype. To document their decisions, they also identified the specific resource at issue (e.g.  time, money, rights, or access to education), selected an indicator associated with the chosen category, highlighted supporting text spans, and provided a brief free-text explanation (1-2 sentences). Where relevant, annotators also marked expressions of the speaker’s own position and references to others’ positions. These additional annotations supported interpretation and documentation during the annotation process, but were not used in the analyses reported in this paper. The full annotation guidelines are provided in Appendix˜G in the Appendix, and Figure˜34 shows an example from the annotation file.
(a) Columns for high-level and (anti-)solidarity categorizations.
(b) Columns for providing explanations.
Figure 34: Example of the annotation process from the annotation file. Figure˜34(a) illustrates the step where annotators choose a high-level label and an (anti-)solidarity subcategory, if applicable. Figure˜34(b) shows columns for detailed explanations, including the choice of a resource, an indicator, and providing free-text commentary
Appendix G Annotation Guidelines Excerpt
This appendix reproduces the annotation guidelines used by the annotators, with minor formatting edits.
General Information
Goal of the study and annotation
The purpose of this study is to examine expressions of solidarity and anti-solidarity towards two specific vulnerable groups—women and migrants—within German parliamentary proceedings. The definitions of (anti-)solidarity are derived from the work Ils et al. (2021), and Eger et al. (2022). We distinguish four subtypes of solidarity and anti-solidarity based on Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of vulnerable groups), exchange-based (contribution, reciprocity, or burden), and empathic (respect for or rejection of diversity).
Details of the Annotation Process
Annotation instructions
In the next section (Annotation Categories), we provide general theoretical information about the categories used in the annotation. This material is intended to support a better understanding of the concepts underlying the task. It is critical that annotators provide a clear explanation for the chosen category or categories in the Explanation column, while keeping comments short. The Other comments column can be used for any additional observations.
Contextual evaluation
Pay close attention to the middle text, which is the sentence whose stance is to be assessed. The surrounding sentences (previous and next) are provided to support contextual interpretation.
Highlighting sentences
When labeling the instances, annotators highlight parts of or full sentences that guided the label choice. Green is used for solidarity categories and red for anti-solidarity categories. In cases of mixed stance, solidarity-expressing parts are highlighted in green and anti-solidarity-expressing parts in red.
Pay attention to the target group
If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None. For example, if the target group is women, then a sentence such as ?all refugees must now leave this country? should also be marked as None.
Structure of the Annotation File
The annotation file contains the following fields: ID: ID of the sentence. The first four numbers after the word debates indicate the year. Category: Indicates the target group (woman or migrant). Pay close attention to the target group. If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None. Sentences: The text to annotate, consisting of Previous, Middle, and Next. The surrounding sentences are provided for contextual interpretation. However, the central focus of the decision should be the middle text. Main category: One of Solidarity, Anti-solidarity, Mixed stance, or None. (Anti-)solidarity subcategory: One of the subcategories, selected only for instances labeled as Solidarity or Anti-solidarity. Explanation: Resource: The resource on the basis of which solidarity or anti-solidarity is expressed (e.g., access to Germany, rights and well-being). Indicators: An indicator motivating the selected anti-solidarity or solidarity type. Indicators are given in drop-down boxes for the respective (anti-)solidarity type and are also listed in the Annotation Categories section. Free text: A short explanation for the chosen label (1–2 sentences). Annotators should provide a clear explanation for the selected category. Other comments: Any additional observations.
Annotation Categories
Definition of solidarity (based on Ils et al. (2021)): solidarity is the readiness to share one’s own resources with others (be it time, financial means, rights, access to education etc.). 1. Solidarity Group-based solidarity Definition. Group-based solidarity stems from a perceived similarity among members of the same group, leading to a shared identity. The group-based solidarity frames often refer to a certain (desired) commonality and a sense of togetherness due to common interests and goals, shared values and norms, or common rights and duties. Characteristics. To identify group-based solidarity, look for: 1. References to a shared identity, common interests, goals, values, or norms within a group (e.g. It’s our collective responsibility to…; we stand together in the fight…; We must unite to…). 2. The existence of mutual obligations or shared rights and duties. 3. Indication that the group must be integrated or even assimilated. 4. Possibly, the identification of an outgroup that contrasts with the in-group. Emphasis on unity and shared goals/duties/rights among the whole group. Compassionate solidarity Definition. Compassionate solidarity is primarily about recognising and supporting individuals or subgroups who are marginalised, disadvantaged, or vulnerable. It extends assistance to these individuals or subgroups, emphasising their need for protection or help. Characteristics. To identify compassionate solidarity, look for: 1. Acknowledgement of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g. We must pay special attention to the needs…; We need to recognize the difficulties…). 2. Emphasis on the need for their protection, assistance, or special consideration due to their marginalised or disadvantaged status (e.g. It is our duty to support…). Emphasis on extra care, support, and protection for those most vulnerable (both in-group and out-group). Exchange-based solidarity Definition. Exchange-based solidarity is linked to the principle of the role of exchange between partners with complementary qualities. Exchange-based solidarity is coded when a speaker refers to the usefulness of ‘exchange partners’ in terms of their actual or future contributions or willingness to contribute. These exchange partners can be rewarded or stimulated, and may also be required to contribute more to receive support. Characteristics. To identify exchange-based solidarity, look for references to a group: 1. That was actively doing, is doing and / or will be doing something useful. 2. Should either be rewarded for past, present or future behavior or that it should be supported to continue that useful behavior. 3. Does not receive enough and needs more support for its contribution (do not confuse it with compassionate solidarity). Empathic solidarity Definition. Empathic solidarity is a concept grounded in the intersubjective valuation of differences among individuals, recognising and respecting their unique characteristics. In contrast to compassionate solidarity, empathic solidarity is coded when a difference is not something that is bad or something that should disappear but is rather a strength or something positive. Empathic solidarity is coded when a speaker acknowledges diversity, difference, or unique characteristics as aspects that should be respected and taken into account. Characteristics. To identify empathic solidarity, look for: 1. Acknowledgement of a group’s diversity and that this diversity must be respected; 2. Recognition of the group’s unique and distinct characteristics; 3. Affirmation of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices. 4. Advocacy for the group’s freedom to develop themselves the way they want. 2. Anti-solidarity Group-based anti-solidarity Definition. The group-based anti-solidarity often points to an out-group characterised by pronounced differences (concerning interests and goals, values and norms, individual rights and responsibilities), and an expression of exclusion and lack of support for that group based on these differences. Characteristics. To identify group-based anti-solidarity, look for: 1. References to differing interests, goals, responsibilities, values, or norms between the groups and at the same time not willing to provide support based on the differences, sense of belonging to different groups. 2. A stronger emphasis on one’s own group’s interests (e.g. It’s every person for themselves…; We must stand up for our own interests…; We must not sacrifice our personal goals for…). 3. Resistance to assimilation or integration with other groups. Compassionate anti-solidarity Definition. Compassionate anti-solidarity is about disregarding and excluding individuals or subgroups either on the basis that they are perceived to be in a “good position” already, or are viewed as undeserving of help. It neglects these individuals or subgroups, downplaying their need for protection or help. Characteristics. To identify compassionate anti-solidarity, look for: 1. Denial of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g. We shouldn’t focus on the problems of…; We don’t need to pay attention to the struggles of…). 2. Minimising the need for their protection, assistance, or special consideration (e.g. It is not our responsibility to aid…) as they are already in a ?good position? or are in a ?bad position?, but still do not deserve help. Exchange-based anti-solidarity Definition. Exchange-based anti-solidarity focuses on the idea that some groups take more than they give. It brings attention to groups that seem to receive a lot of resources or support, but do not contribute much or cause problems. This perspective supports punishing such groups, giving them less help, or making them contribute more. Characteristics. To identify exchange-based anti-solidarity, look for references to a group: 1. Talk about a group that has received a lot, but has not given much back or has caused harm. 2. Suggestions that this group should be punished, or get less than they are currently getting. 3. Calls for the group to give or do more in return for help. Empathic anti-solidarity Definition. Empathic anti-solidarity refers to the cases when a speaker refers to a group indicating that this group is different from everyone else, but this difference (needs, characteristics, beliefs, opinions) should not be respected or recognised. The speaker could argue that this difference should not be acknowledged. The core logic behind this is: ?You’re not like me, and that’s a shame. I don’t understand or respect you or your differences?. Characteristics. To identify empathic anti-solidarity, look for: 1. Disregard for a group’s diversity, arguing that this diversity should not be respected. 2. Ignorance or dismissal of the group’s unique and distinct characteristics. 3. Denial of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices. 3. Mixed stance Definition. Mixed stance refers to instances where the speaker expresses both support and opposition towards the target group within the same discourse or text. It involves contradictions or conflicting statements regarding the speaker’s stance towards the target group. Characteristics. 1. Annotate instances where the speaker presents a mixed stance by expressing both support and opposition towards the target group in the same discourse or text. 2. Look for contradictions or inconsistencies in the level of support or opposition expressed towards the target group. 3. Note that the middle sentence has more weight. 4. None Definition. None refers to instances where the speaker neither expresses explicit support nor opposition towards the target group. The statement is unrelated to either solidarity or anti-solidarity or is unrelated to the group altogether. Characteristics. 1. Annotate instances where the speaker maintains a neutral position and does not express clear support or opposition towards the target group. 2. Look for statements that provide factual information, or refrain from taking a stance or expressing personal opinion. 3. Also label the instance as None if you think that it is unrelated to the target group.